A firm produces candles.The market for candles is highly competitive,with candles currently selling for $10.The firm's shortrun total cost function is C=200+0.2q^2 ,so it's marginal cost is MC=0.4
A) is the firm earning a profit?
B) what is the shortrun shutdown price?
1
Expert's answer
2017-04-12T10:43:05-0400
a)MC=P

0.4*Q=10

Q=25

Profit=TR-TC=25*10- 200-0.2*25^2=250- 200-125=- 75

b)VC=125, AVC=5- shortrun shutdown price